OpenStudy (anonymous):

Simplify Completely? Multiple choice ((7x^2)+11x-6)/((7^2)-10x+3)

OpenStudy (anonymous):

\[\frac{ 7x^2+11x-6 }{ 7x^2-10x+3 }\]

OpenStudy (anonymous):

\[\frac{ x+2 }{ 7x+3 }\]

OpenStudy (anonymous):

HINT: \[7x ^{2}+11x-6 = 7x ^{2} + 14x - 3x - 6\]

OpenStudy (anonymous):

\[\frac{ x+2 }{ x-1 }\]

OpenStudy (anonymous):

\[\frac{ 7x-6 }{ x-1 }\]

OpenStudy (anonymous):

\[7x ^{2} - 10x + 3 = 7x ^{2} - 7x - 3x + 3\]

OpenStudy (anonymous):

\[\frac{ 7x-6 }{ 7x+3 }\]

OpenStudy (anonymous):

\[\frac{ (7x-3)(x+2) }{ (7x-3)(x-1) }\]

OpenStudy (anonymous):

Foil? @Soumyadeep

OpenStudy (anonymous):

\[\frac{ x+2 }{ x-1 }\]

OpenStudy (anonymous):

I got it! Thanks for the help guys!